GreenHab report – December 14th

Crew 200 Greenhab Report 14-DEC-2018

Greenhab Officer:
Makiah Eustice

Environmental Control:
40% Shade Cloth on

Average Temperatures:
Low: 24 C
High: 50 C

Hours of Supplemental Light:

Daily Water Usage of Crops:
8 gallons

Water in Blue Tank (lbs):

Times of Watering for Crops:

Changes to crops:

The high measurement was still 50 today, but since I wasn’t at the
greenhab at the highest time in the day.

1.6 g of Sage
95.8 g Sunflower Sprouts
37.8 g Mixed Salad Greens
4.4 g Cilantro
6.4 g Water Cress

Support/Supplies Needed:
Was there suppose to be a weight in the GreenHab? If so, I cannot find
it and have to go to the Science Dome to get the mass of the harvest.
GreenHab is a great place to work, it has alot of potential. With
added sensors and automated suns shades, the Hab GreenHab can become
closer to the type of system we’d want on Mars!

Sol Summary – December 14

Crew 200 Sol Summary Report 14-DEC-2018

Sol: 12

Summary Title: Driving to the moons and cleaning the habitat

Author’s name: Lindsay Rutter

Mission Status: Simulation ended at 1:00PM today. All crew members
happy and healthy.

Sol Activity Summary: We spent much of the morning finishing our
reports about this week. Each day, we have been documenting our
observations of the dynamics of a 3 member crew. Today, we summarized
it over a decadent dessert. One crew member made cinnamon chocolate
cake by modifying “Kyle’s Carrot Cake”, a recipe in the habitat
cookbook “Mars Home Cooking: Recipes from the Mars Desert Research
Station”. After we discontinued our simulation as planned, we drove
two rovers to an array of “moons” north of the habitat. White Moon.
Beige Moon. Yellow Moon. Gray Moon.

Look Ahead Plan: Tonight and tomorrow, we will be thoroughly cleaning
the habitat and its surrounding buildings. Tomorrow, we will also
drive to Hanksville to pick up the next crew and drive them to MDRS.
We will introduce the MDRS to them and have discussions with them
about the rules and expectations of conducting missions here.

Anomalies in work: None reported

Weather: Snow melting; chilly with moderate cloud coverage.

Crew Physical Status: No crew members reported physical problems today.


Reports to be filed: Operations Report, GreenHab Report, Pictures
Report, Journalist Report, Sol Report, Mission Summary Report

Support Requested: N/A

Operations Report – December 14th

Crew 200 Operations Report 14-Dec-2018

SOL: 12

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 4:50 pm last night to 9 am today. On tonight at 5:00pm.

Solar— SOC 80%
Diesel Reading – 97 % (refilled)
Propane Reading – 26 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 330 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 4

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: no
Hours: 43
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: yes
Hours: 62.6
Beginning charge: 100%
Ending charge: 31%
Currently charging: yes

Opportunity rover used: yes
Hours: 43.4
Beginning charge: 100%
Ending charge: 35%
Currently charging: yes

Curiosity rover used: no
Hours: 58.8
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: we used the rover for an out-of-sim drive to the north
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: after working on our projects this morning, we left sim at noon, spent the afternoon riding/hiking in the north and came back at 4pm to clean the science dome floor, Hab floor and the RAM.

Summary of internet: we used more than the daily quantity today, this was mainly due to our projects.

Summary of suits and radios: Nothing to report, spacesuit 3 does not have a functional ventilation. We received an email for this, I will try to have a look tomorrow if I have the time.

Summary of Hab operations: we began cleaning.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Floor mopped and cleaned again, looks better.

Summary of RAM operations: Cleaned up.

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site